\begin{thebibliography}{31}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Ahn and Mortensen(2019)}]{ahn2019}
Emily Ahn and David Mortensen. 2019.
\newblock \href {https://typology-and-nlp.github.io/2019/proceedings.pdf}
  {{Predicting Continuous Vowel Spaces in the Wilderness}}.
\newblock In \emph{Proc. of TyP-NLP: The First Workshop on Typology for
  Polyglot NLP}, pages 22--24, Florence, Italy. Association for Computational
  Linguistics.

\bibitem[{Anastasopoulos(2019)}]{anastasopoulos2019}
Antonios Anastasopoulos. 2019.
\newblock \emph{Computational Tools for Endangered Language Documentation}.
\newblock Ph.D. thesis, University of Notre Dame, Indiana.

\bibitem[{Baljekar et~al.(2015)Baljekar, Sitaram, Muthukumar, and
  Black}]{baljekar2015}
Pallavi Baljekar, Sunayana Sitaram, Prasanna~Kumar Muthukumar, and Alan~W
  Black. 2015.
\newblock \href
  {https://www.isca-speech.org/archive/interspeech_2015/i15_3194.html} {Using
  articulatory features and inferred phonological segments in zero resource
  speech processing}.
\newblock In \emph{Proc. Interspeech}, pages 3194--3198, Dresden, Germany.
  ISCA.

\bibitem[{Chen et~al.(2017)Chen, Hasegawa-Johnson, Chen, and Lim}]{chen2017}
Wenda Chen, Mark Hasegawa-Johnson, Nancy~F. Chen, and Boon~Pang Lim. 2017.
\newblock \href {http://dx.doi.org/10.21437/Interspeech.2017-1567} {{Mismatched
  Crowdsourcing from Multiple Annotator Languages for Recognizing
  Zero-Resourced Languages: A Nullspace Clustering Approach}}.
\newblock In \emph{Proc. Interspeech}, pages 2789--2793, Stockholm, Sweden.
  ISCA.

\bibitem[{Chen et~al.(2019)Chen, Tu, chieh Yeh, and Lee}]{chen2019}
Yuan-Jui Chen, Tao Tu, Cheng chieh Yeh, and Hung-Yi Lee. 2019.
\newblock \href {http://dx.doi.org/10.21437/Interspeech.2019-2730} {{End-to-End
  Text-to-Speech for Low-Resource Languages by Cross-Lingual Transfer
  Learning}}.
\newblock In \emph{Proc. Interspeech}, pages 2075--2079, Graz, Austria. ISCA.

\bibitem[{Cotterell and Eisner(2017)}]{cotterell-eisner-2017-probabilistic}
Ryan Cotterell and Jason Eisner. 2017.
\newblock \href {https://doi.org/10.18653/v1/P17-1109} {Probabilistic typology:
  Deep generative models of vowel inventories}.
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1182--1192,
  Vancouver, Canada. Association for Computational Linguistics.

\bibitem[{Dalsgaard and Andersen(1992)}]{dalsgaard1992}
Paul Dalsgaard and Ove Andersen. 1992.
\newblock \href {https://www.isca-speech.org/archive/icslp_1992/i92_0547.html}
  {{Identification of Mono- and Poly-Phonemes Using Acoustic-Phonetic Features
  Derived by a Self-Organising Neural Network}}.
\newblock In \emph{Proc. 2nd International Conference on Spoken Language
  Processing (ICSLP)}, pages 547--550, Banff, Canada.

\bibitem[{Dryer and Haspelmath(2013)}]{wals}
Matthew~S. Dryer and Martin Haspelmath. 2013.
\newblock {WALS} online.
\newblock \emph{Leipzig: Max Planck Institute for Evolutionary Anthropology}.
\newblock \url{http://wals.info/}.

\bibitem[{Frankel et~al.(2007)Frankel, Wester, and King}]{frankel2007}
Joe Frankel, Mirjam Wester, and Simon King. 2007.
\newblock \href {https://doi.org/10.1016/j.csl.2007.03.002} {Articulatory
  feature recognition using dynamic {B}ayesian networks}.
\newblock \emph{Computer Speech \& Language}, 21(4):620--640.

\bibitem[{Gordon(2016)}]{gordon2016}
Matthew~K Gordon. 2016.
\newblock \emph{Phonological Typology}.
\newblock Oxford Surveys in Phonology and Phonetics. Oxford University Press.

\bibitem[{Gutkin(2020)}]{gutkin2020}
Alexander Gutkin. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.sltu-1.2} {{E}idos: An
  open-source auditory periphery modeling toolkit and evaluation of
  cross-lingual phonemic contrasts}.
\newblock In \emph{Proc. 1st Joint Workshop on Spoken Language Technologies for
  Under-resourced languages (SLTU) and Collaboration and Computing for
  Under-Resourced Languages (CCURL)}, pages 9--20, Marseille, France. ELRA.

\bibitem[{Gutkin et~al.(2018)Gutkin, Merkulova, and Jansche}]{gutkin2018}
Alexander Gutkin, Tatiana Merkulova, and Martin Jansche. 2018.
\newblock \href {http://dx.doi.org/10.21437/SLTU.2018-52} {{Predicting the
  Features of {W}orld {A}tlas of {L}anguage {S}tructures from Speech}}.
\newblock In \emph{Proc. 6th International Workshop on Spoken Language
  Technologies for Under-Resourced Languages (SLTU)}, pages 248--252, Gurugram,
  India. ISCA.

\bibitem[{Hu et~al.(2020)Hu, Settle, and Livescu}]{hu2020}
Yushi Hu, Shane Settle, and Karen Livescu. 2020.
\newblock \href {https://arxiv.org/pdf/2006.14007.pdf} {Multilingual jointly
  trained acoustic and written word embeddings}.
\newblock \emph{arXiv preprint arXiv:2006.14007}.

\bibitem[{Jansche and Gutkin(2019)}]{jansche2019}
Martin Jansche and Alexander Gutkin. 2019.
\newblock \href {http://dx.doi.org/10.21437/Interspeech.2019-2740} {{Sampling
  from Stochastic Finite Automata with Applications to CTC Decoding}}.
\newblock In \emph{Proc. Interspeech}, pages 2230--2234, Graz, Austria. ISCA.

\bibitem[{Johny et~al.(2019)Johny, Gutkin, and Jansche}]{johny2019}
Cibu Johny, Alexander Gutkin, and Martin Jansche. 2019.
\newblock \href {http://dx.doi.org/10.21437/Interspeech.2019-2184}
  {{Cross-Lingual Consistency of Phonological Features: An Empirical Study}}.
\newblock In \emph{Proc. Interspeech}, pages 1741--1745, Graz, Austria. ISCA.

\bibitem[{Lee et~al.(2015)Lee, O'Donnell, and Glass}]{lee2015}
Chia-ying Lee, Timothy~J O'Donnell, and James Glass. 2015.
\newblock \href {https://doi.org/10.1162/tacl_a_00146} {Unsupervised lexicon
  discovery from acoustic input}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  3:389--403.

\bibitem[{Li et~al.(2020)Li, Dalmia, Li, Littell, Lee, Yao, Anastasopoulos,
  Mortensen, Neubig, Black, and Metze}]{li2020}
Xinjian Li, Siddharth Dalmia, Juncheng Li, Patrick Littell, Matthew Lee, Jiali
  Yao, Antonios Anastasopoulos, David Mortensen, Graham Neubig, Alan Black, and
  Florian Metze. 2020.
\newblock \href {https://doi.org/10.1109/ICASSP40776.2020.9054362} {{Universal
  Phone Recognition with a Multilingual Allophone System}}.
\newblock In \emph{Proc. 45th International Conference on Acoustics, Speech,
  and Signal Processing (ICASSP)}, pages 8249--8253, Barcelona, Spain. IEEE.

\bibitem[{Littell et~al.(2017)Littell, Mortensen, Lin, Kairis, Turner, and
  Levin}]{littell-etal-2017-uriel}
Patrick Littell, David~R. Mortensen, Ke~Lin, Katherine Kairis, Carlisle Turner,
  and Lori Levin. 2017.
\newblock \href {https://www.aclweb.org/anthology/E17-2002} {{URIEL} and
  lang2vec: Representing languages as typological, geographical, and
  phylogenetic vectors}.
\newblock In \emph{Proceedings of the 15th Conference of the {E}uropean Chapter
  of the Association for Computational Linguistics: Volume 2, Short Papers},
  pages 8--14, Valencia, Spain. Association for Computational Linguistics.

\bibitem[{Moran and McCloy(2019{\natexlab{a}})}]{phoible380}
Steven Moran and Daniel McCloy. 2019{\natexlab{a}}.
\newblock \href {https://phoible.org/inventories/view/380} {{Javanese sound
  inventory from UCLA Phonological Segment Inventory Database (UPSID)}}.
\newblock In \emph{PHOIBLE 2.0}. Max Planck Institute for the Science of Human
  History, Jena.

\bibitem[{Moran and McCloy(2019{\natexlab{b}})}]{phoible}
Steven Moran and Daniel McCloy, editors. 2019{\natexlab{b}}.
\newblock \href {https://phoible.org/} {\emph{PHOIBLE 2.0}}.
\newblock Max Planck Institute for the Science of Human History, Jena.

\bibitem[{Morita and Koda(2020)}]{morita2020}
Takashi Morita and Hiroki Koda. 2020.
\newblock \href {https://arxiv.org/abs/2005.05487} {Exploring {TTS} without {T}
  using {B}iologically/{P}sychologically {M}otivated {N}eural {N}etwork
  {M}odules ({Z}ero{S}peech 2020)}.
\newblock \emph{arXiv preprint arXiv:2005.05487}.

\bibitem[{Moritz et~al.(2019)Moritz, Hori, and Roux}]{moritz2019}
N.~Moritz, T.~Hori, and J.~L. Roux. 2019.
\newblock \href {http://dx.doi.org/10.1109/ICASSP.2019.8683510} {{Triggered
  Attention for End-to-end Speech Recognition}}.
\newblock In \emph{Proc. 44th International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 5666--5670, Brighton, United Kingdom.
  IEEE.

\bibitem[{Mortensen et~al.(2020)Mortensen, Li, Littell, Michaud, Rijhwani,
  Anastasopoulos, Black, Metze, and Neubig}]{mortensen2020}
David~R. Mortensen, Xinjian Li, Patrick Littell, Alexis Michaud, Shruti
  Rijhwani, Antonios Anastasopoulos, Alan~W. Black, Florian Metze, and Graham
  Neubig. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.lrec-1.656/} {{AlloVera:
  A Multilingual Allophone Database}}.
\newblock In \emph{Proc. 12th Language Resources and Evaluation Conference
  (LREC)}, pages 5329--5336, Marseille, France. ELRA.

\bibitem[{Mortensen et~al.(2016)Mortensen, Littell, Bharadwaj, Goyal, Dyer, and
  Levin}]{mortensen-etal-2016-panphon}
David~R. Mortensen, Patrick Littell, Akash Bharadwaj, Kartik Goyal, Chris Dyer,
  and Lori Levin. 2016.
\newblock \href {https://www.aclweb.org/anthology/C16-1328} {{P}an{P}hon: A
  resource for mapping {IPA} segments to articulatory feature vectors}.
\newblock In \emph{Proceedings of {COLING} 2016, the 26th International
  Conference on Computational Linguistics: Technical Papers}, pages 3475--3484,
  Osaka, Japan. The COLING 2016 Organizing Committee.

\bibitem[{M{\"u}ller et~al.(2017)M{\"u}ller, Franke, Waibel, and
  St{\"u}ker}]{muller2017}
Markus M{\"u}ller, J{\"o}rg Franke, Alex Waibel, and Sebastian St{\"u}ker.
  2017.
\newblock \href {https://doi.org/10.1109/ICASSP.2017.7953148} {Towards phoneme
  inventory discovery for documentation of unwritten languages}.
\newblock In \emph{Proc. 42nd International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 5200--5204, New Orleans, USA. IEEE.

\bibitem[{van Niekerk et~al.(2020)van Niekerk, Nortje, and
  Kamper}]{niekerk2020}
Benjamin van Niekerk, Leanne Nortje, and Herman Kamper. 2020.
\newblock \href {https://arxiv.org/abs/2005.09409} {Vector-quantized neural
  networks for acoustic unit discovery in the {Z}ero{S}peech 2020 challenge}.
\newblock \emph{arXiv preprint arXiv:2005.09409}.

\bibitem[{Ogloblin(2006)}]{ogloblin2006}
Alexander Ogloblin. 2006.
\newblock \href
  {https://www.routledgehandbooks.com/doi/10.4324/9780203821121.ch1}
  {Javanese}.
\newblock In Alexander Adelaar and Niklaus Himmelman, editors, \emph{The
  Austronesian Languages of Asia and Madagascar}, 1st edition, chapter~21,
  pages 590--–594. Routledge.

\bibitem[{Ondel et~al.(2019)Ondel, Li, Sell, and Hermansky}]{ondel2019}
Lucas Ondel, Ruizhi Li, Gregory Sell, and Hynek Hermansky. 2019.
\newblock \href {https://doi.org/10.1109/ICASSP.2019.8682787} {{Deriving
  Spectro-Temporal Properties of Hearing from Speech Data}}.
\newblock In \emph{Proc. 44th International Conference on Acoustics, Speech and
  Signal Processing (ICASSP)}, pages 411--415, Brighton, UK. IEEE.

\bibitem[{Salesky and Black(2020)}]{salesky2020}
Elizabeth Salesky and Alan~W Black. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.acl-main.217} {{Phone
  Features Improve Speech Translation}}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 2388--2397. ACL.

\bibitem[{Skidmore and Gutkin(2020)}]{skidmore2020}
Lucy Skidmore and Alexander Gutkin. 2020.
\newblock \href {https://doi.org/10.1007/978-3-030-60276-5_51} {{Does A Priori
  Phonological Knowledge Improve Cross-Lingual Robustness of Phonemic
  Contrasts?}}
\newblock In \emph{Proc. 22nd International Conference on Speech and Computer
  (SPECOM)}, pages 530--543, St. Petersburg, Russia. Springer Nature.

\bibitem[{Tjandra et~al.(2019)Tjandra, Sakti, and Nakamura}]{tjandra2019}
Andros Tjandra, Sakriani Sakti, and Satoshi Nakamura. 2019.
\newblock \href {https://doi.org/10.1109/ASRU46091.2019.9003853}
  {{Speech-to-Speech Translation between Untranscribed Unknown Languages}}.
\newblock In \emph{Proc. Automatic Speech Recognition and Understanding
  Workshop (ASRU)}, pages 593--600. IEEE.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{emnlp2020}
\usepackage[utf8]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{times}
\usepackage{latexsym}
\usepackage[flushmargin]{footmisc}

\usepackage{microtype}

\aclfinalcopy 
\def\aclpaperid{SIGTYP 5}

\title{Towards Induction of Structured Phoneme Inventories}

\author{Alexander Gutkin\\
  Google, UK\\
  {\small\texttt{agutkin@google.com}}\\\And
  Martin Jansche\thanks{\ \ This work was done while the author was at Google, prior to joining Amazon.}\\
  Amazon, UK\\
  {\small\texttt{jansche@amazon.com}}\\\And
  Lucy Skidmore\\
  University of Sheffield, UK\\
  {\small\texttt{lskidmore1@sheffield.ac.uk}}\\}

\date{}

\begin{document}
\maketitle

\vspace{-1.5mm}
\section{Introduction}
\vspace{-1.7mm}
Phonological typology is an important branch of linguistic typology
concerned with the study of the distribution and behavior of sounds in
world's languages~\cite{gordon2016}. The cross-linguistic typological
databases, such as \textsc{phoible} by~\citet{phoible}, provide the crucial tools
for drawing typological generalizations within the field. In addition
to facilitating development of probabilistic models of phonological
typology~\cite{cotterell-eisner-2017-probabilistic,ahn2019}, such
resources were also shown to positively influence the downstream
multilingual NLP~\cite{littell-etal-2017-uriel}, speech~\cite{li2020}
and language documentation tasks~\cite{anastasopoulos2019}.

This short paper provides an overview of our completed and ongoing
experiments with phonological representations utilizing phonological
typology databases in speech processing tasks. Despite the increasing
popularity of end-to-end approaches to automatic speech
recognition~\cite{moritz2019}, text-to-speech~\cite{chen2019} and
speech-to-speech translation~\cite{tjandra2019}, there are still
plenty of scenarios where integration of accurate phonological
knowledge is crucial, or at least beneficial, including the end-to-end
approaches themselves as recently demonstrated by~\citet{salesky2020}.

In particular, we describe two strains of research motivated by the
need to scale the development of speech technologies that still
require phonological representations to low-resource languages and
dialects. We first overview the framework for analysis of
cross-lingual consistency of phonological features in multilingual
phoneme inventories derived from cross-lingual typological
databases. We then offer a sketch of a method that may serve as a
potential building block in the future phoneme inventory induction
system and the central role the phonological typology plays in this
approach.

\vspace{-1mm}
\section{Multilingual Phoneme Inventories}
\vspace{-1.7mm}
The phoneme was originally defined as a theoretical abstraction that
applies language-internally. Using phonemes and their succinct
distinctive feature (DF) encodings in cross-linguistic settings, the
practice going back to~\citet{dalsgaard1992}, raises an important
question: given a multilingual phoneme inventory derived from a
typological database, such as \textsc{phoible} or
\textsc{PanPhon}~\cite{mortensen-etal-2016-panphon}, it is not clear a priori
whether all the DFs will be useful or even valid. If DF
representations were phonetic rather than phonemic, and acoustic
rather than articulatory, one would expect a close correspondence
between DFs and the acoustic signal. In practical multilingual
applications, however, the representations are often guided by purely
phonemic considerations because of the availability of phonemic
inventories and transcriptions. Such approaches are often outperformed
by the models striving for more phonetic realism, such as the
allophonic models of~\citet{mortensen2020}.

We followed a simple method: to consider a phonemic contrast to be
consistent or robust across languages, it needs to be easily predictable
on heldout languages in a binary classification
task~\cite{johny2019}. An instance of this problem consists of a span
of a speech signal (e.g., a vowel in surrounding context) and a
positive or negative label (e.g., front vowel vs. back vowel). A
classifier is trained on a multi-speaker, multi-language dataset
withholding one or more languages, which are later used for
evaluation. For cases where cross-linguistic consistency did not
hold, we extended this method by additionally grounding the
representation on the contextual phonological knowledge given as DFs,
excluding the contrast itself~\cite{skidmore2020}. For our experiments
we used a set of languages from Dravidian, Indo-European
and Malayo-Polynesian families with the phoneme inventories derived
from \textsc{phoible}.

Overall, our findings are mixed. A specially designed experiment for
predicting contrasts in unvoiced labial consonants between Bengali and
Spanish produced consistent and cross-lingually robust
predictions~\cite{johny2019}, also for a variety of auditory
representations~\cite{gutkin2020}, despite the conflicting status of
some of the allophones of the phonemes in the experiment. Similarly
robust were contrasts between front and back vowels, as well as the
vowel height and continuant manner of articulation
distinctions~\cite{skidmore2020}. The negative results include the
cross-lingual prediction of retroflex consonants between the language
families: retroflex predictor trained on Dravidian languages fails to
reliably predict retroflex consonants in Bengali, conversely the
predictor trained on Indo-Aryan languages is not reliable for
Malayalam. Similarly inconsistent, but less disappointing, was the
detection of aspiration. Furthermore, inclusion of other contrasts as
contextual input features did not lead to significant improvements in
predicting these hard cases.

One of the motivations behind our investigations describe above is a
research question that still remains open: Can the above methodology
be used for analyzing the cross-linguistic quality of the existing
phoneme inventories given the data? For example, among the
Malayo-Polynesian languages we considered, only Javanese has
retroflex consonants, which it acquired through loanwords from
Indo-Aryan or Dravidian languages~\cite{ogloblin2006}. \textsc{phoible}
contains three different phoneme inventories for Javanese, the largest
of which \texttt{GM} \texttt{1675} represents retroflex plosives,
while the smallest \texttt{UPSID} \texttt{380}~\cite{phoible380} omits
them. Which of the two representations is more likely to be a better
fit in a multilingual pronunciation model?

\vspace{-1mm}
\section{Towards Phonology Induction}
\vspace{-1.7mm}
Our ongoing research focuses on induction of phoneme inventories for
languages for which the conventional resources required for speech
model training are often missing, placing this task among other
related work in zero-resource subword modeling
problems~\cite{baljekar2015,lee2015,chen2017}. The current
state-of-the-art unsupervised acoustic unit discovery approaches
derive acoustic-phonetic~\cite{hu2020,morita2020,niekerk2020} and
latent auditory-like~\cite{ondel2019} representations, yet it is
unclear how accurate these representations are from a typological
standpoint.

Our initial investigations utilized ``universal'' multilingual phoneme
recognizers. The results were disappointing primarily because the
training data sparsity and the absence of language models to constrain
the search often resulted even in unreliable recovery of the phoneme
inventories for unseen dialects of the same language, e.g.,
determining the inventory of Argentinian Spanish having seen Castilian
and Mexican Spanish. A more viable approach to this task is
integrating language identification and phonological typology into the
phoneme recognizer. Given an accurate language identification model,
the \textsc{phoible} phoneme inventories belonging to the ``closest''
languages may be used to constrain the phonemic hypothesis space for
the unseen language or dialect. We developed efficient search
techniques for the lattices of this type~\cite{jansche2019}.

Currently we are revisiting the phonological contrast predictor
methods described in the previous section to adapt them to phonology
induction tasks. Several known approaches to phonemic feature
detection in continuous speech are known, some are purely reliant on
signal processing, while others are
model-based~\cite{frankel2007,muller2017}. At the simplest level, the
output of such predictors represents speech as the parallel
asynchronous streams of articulatory features. More sophisticated
models that exploit the structure of articulation, feature geometry
and other types of correlations between various features are also
possible. In these approaches, the phonological cross-linguistic
databases, such as \textsc{phoible} or \textsc{PanPhon}, provide the important source
of truth not only for combining various features into phonemes, but
also for determining which combinations of hypothesized phonemes are
admissible given the existing phoneme inventories. Furthermore,
additional phonological evidence provided by other typological
resources, such as World Atlas of Language Structures
(\textsc{WALS})~\cite{wals}, can be integrated as well if it can be
reliably extracted from the speech signal~\cite{gutkin2018}.

\bibliographystyle{acl_natbib}
\bibliography{anthology,main}

\end{document}